# Anytime-valid t-tests and confidence sequences for Gaussian means with unknown variance

### v3

@@ 1 Introduction
Lai and more recent authors [Grünwald et al., 2023, Pérez-Ortiz et al., 2022] situate the Gaussian t-test problem into the broader framework of group-invariant tests; indeed, the t-test null is invariant under rescaling of observations by a constant. The tests are mostly constructed using variants of the sequential likelihood ratio test (LRT) by Wald [1945], utilizing the fact that the likelihood ratios are nonnegative martingales (see Section 2.5). Building upon these group invariance tools, Lindon et al. [2022] construct a sequential F-test for the regression set-up, which coincides with one of the t-tests that we shall present when no covariate exists. A slightly different approach is taken by Jourdan et al. [2023] who study the same problem in the bandit context, sequentially bounding the empirical mean and variance separately using existing Z-test-type bounds and obtaining via a union bound a sequence of confidence “boxes” for mean and variance simultaneously. We shall return to and compare our bounds with some of these results later.
@@
    We derive the information-theoretic lower bound on the width of confidence intervals, and upper bound on the growth of e-processes for the t-test, both of which are attained, or almost attained, by our results (Contributions 1 and 3 in this list). We also show that other conservative sequential t-tests available in the literature (Lai’s original method, the plug-in method by Jourdan et al. [2023], the median method by Howard and Ramdas [2022]), and very surprisingly, even the fixed-time classical t-test, fail to be optimal in these regards.
@@ 2.2 Sequential Statistics
2.3 Test Supermartingales and Extensions
@@
Lemmas 2.1 and 2.2 suggest that, if $\mathcal{P}_{0}$ is a set of null distributions, NSMs, e-processes, and ENSMs for $\mathcal{P}_{0}$ can be seen as measurements of evidence against the null $H_{0}:P\in\mathcal{P}_{0}$. To see that, if we have at hand an NSM or e-process $\{M_{n}\}$ for $\mathcal{P}_{0}$ such that $M_{0}=1$, we may reject the null $\mathcal{P}_{0}$ whenever the process $M_{n}$ exceeds $1/\alpha$. This, due to Lemma 2.1, controls the sequential type 1 error $P(H_{0}\text{ is ever rejected})$ within $\alpha$ for any $P\in\mathcal{P}_{0}$. Similarly, if we have an ENSM $\{M_{n}\}$ for $\mathcal{P}_{0}$, we may set the right hand side of (5) to $\alpha$ and solve the corresponding $\varepsilon$, rejecting $H_{0}:P\in\mathcal{P}_{0}$ whenver $M_{n}$ exceeds $\varepsilon$. Desirably, an NSM, e-process or ENSM for $\mathcal{P}_{\theta_{0}}$ used as a powerful test process shrinks under the null $\mathcal{P}_{0}$ and grows under the alternative $\mathcal{P}_{1}$. A confidence sequence for a parameter $\theta$ follows in the same way if we have an NSM or an ENSM $\{M^{\theta_{0}}_{n}\}$ for every $\mathcal{P}_{\theta_{0}}=\{P\in\mathcal{P}:\theta(P)=\theta_{0}\}$: by solving the inequality $M_{n}^{\theta_{0}}\leqslant 1/\alpha$ or $M_{n}^{\theta_{0}}\leqslant\varepsilon$ in $\theta_{0}$, that is. An illustrative comparison between NSMs and ENSMs in the Gaussian case with known variance for both constructing tests and confidence sequences is provided recently by Wang and Ramdas [2023a, Section 5].
@@ 2.5 Likelihood Ratio Martingales
As a test process, the e-process $\{R_{n}\}$ distinguishes the null $\mathcal{N}_{\mu=0}$ and the alternative $\mathcal{N}_{\mu\neq 0}$ due to the following asymptotic result on a limit which we refer to as the “growth index” of the test process.
Proposition 3.3 (Growth index of the universal inference t-test e-process).
@@
The $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ growth index in Proposition 3.3, we remark, is universal among test processes for t-tests. If one uses the plug-in likelihood ratio $\{\ell_{n}^{0,\sigma^{2}}\}$ in Corollary 3.1 to conduct the Z-test for the null $N_{0,\sigma^{2}}$, then, under the actual distribution $N_{\mu,\sigma^{2}}$, it is not hard to see that a faster rate $\frac{\log\ell_{n}^{0,\sigma^{2}}}{n}\to\mu^{2}/2\sigma^{2}$ holds under similar assumptions of the point estimators (which we show in Appendix B). We shall see in the rest of the paper many more occurrences of the same $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ limit, and we shall discuss that this is indeed the optimal rate in Section 6.3.
@@
Proposition 3.5 (Growth index of the universal inference one-sided t-test e-process).
@@ 4.4 An Extended Test Martingale for t-Test
$\{H_{n}\}$ works favorably as a test process for the null $\mathcal{N}_{\mu=0}$ and the alternative $\mathcal{N}_{\mu\neq 0}$ due to the following symptotic result on both its growth index under the alternative and its convergence under the null.
@@ 5.1 Theoretical Comparison
Growth index $\mu^{2}/2\sigma^{2}$ $(\mu\vee 0)^{2}/2\sigma^{2}$ $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ $\frac{1}{2}\log(1+(\mu\vee 0)^{2}/\sigma^{2})$
Table 2: Comparison of universal inference-based Z-tests and t-tests, where $K_{n}:=\prod_{i=1}^{n}\frac{\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}}{\widetilde{\sigma}_{i-1}}$. The “growth index” of a process $M_{n}$ refers to the almost sure limit of $\frac{\log M_{n}}{n}$ under the true distribution $N_{\mu,\sigma^{2}}$.
@@
Growth index $\mu^{2}/2\sigma^{2}$ $(\mu\vee 0)^{2}/2\sigma^{2}$ $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ $\frac{1}{2}\log(1+(\mu\vee 0)^{2}/\sigma^{2})$
@@ 6 Optimality
In this section, we show that the $\mathcal{O}(\alpha^{-1/n})$ or $\mathcal{O}(\alpha^{-1/(n-1)})$ growth rate (when $n$ is fixed and $\alpha\to 0$) in our t-confidence sequences shown in Table 1, and that the $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ growth index in our t-test processes shown in Tables 2 and 3, are both optimal. These two kinds of optimality require different techniques, but the Kullback-Leibler divergence
@@ 6.2 The Surprising $\alpha$-Suboptimality of the Classical t-Test
To see that this is an e-value without invoking the full Theorem 3.2, use the fact that $|x-\mu|\leqslant\sigma\exp((x-\mu)^{2}/2\sigma^{2}+1/2)$ for all $x,\sigma$ to conclude that
@@
$\mathrm{E}_{N_{\mu,\sigma^{2}}}\left(\sigma\exp\left(\frac{(X_{1}-\mu)^{2}}{2\sigma^{2}}+\frac{1}{2}\right)\cdot\exp\left(\frac{1-X_{1}^{2}}{2}\right)\right)=\mathrm{E}_{N_{\mu,\sigma^{2}}}\left(\frac{\mathrm{d}N_{0,1}}{\mathrm{d}N_{\mu,\sigma^{2}}}(X_{1})\right)=1.$ (59)
@@
6.3 Reverse Information Projection and Numeraire
We now show that the recurring growth index $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ in various test processes presented thus far is not a coincidence, as it is the optimal rate of growth in the problem of t-test.
@@ 6.3 Reverse Information Projection and Numeraire
It is known that when testing the point null $P$, if the real distribution $Q\neq P$, an e-process $\{M_{n}\}$ grows optimally at the rate of
$\frac{\mathrm{E}_{Q}(\log M_{n})}{n}\leqslant D_{\operatorname{KL}}(Q\|P),$ (62)
@@
with equality if and only if $\{M_{n}\}$ is the likelihood ratio in Lemma 2.3. See e.g. Vovk and Wang [2022, Lemma 2.1]. For example, in sequential Z-tests, $P=N_{0,\sigma^{2}}$ and $Q=N_{\mu,\sigma^{2}}$, then the upper bound reads
$D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}\|N_{0,\sigma^{2}})=\frac{\mu^{2}}{2\sigma^{2}}.$ (63)
@@
More generally, if we are to test a composite null $\mathcal{P}$ and the real distribution $Q\notin\mathcal{P}$, an e-process $\{M_{n}\}$ grows at most at the rate of
$\frac{\mathrm{E}_{Q}(\log M_{n})}{n}\leqslant\inf_{P\in\mathcal{P}}D_{\operatorname{KL}}(Q\|P).$ (64)
If the infimum $\inf_{P\in\mathcal{P}}D_{\operatorname{KL}}(Q\|P)$ is attained as minimum by some $P^{*}\in\mathcal{P}$, we say that $P^{*}$ is the reverse information projection of $Q$ onto the set $\mathcal{P}$; if, further, the likelihood ratio process $M_{n}^{*}=\prod_{i=1}^{n}\frac{\mathrm{d}Q}{\mathrm{d}P^{*}}(X_{i})$ is an e-process for $\mathcal{P}$, then $\{M_{n}^{*}\}$ is the optimal (fastest growing) e-process for $\mathcal{P}$ under the real distribution $Q$, the one that attains equality in (64), and is called the numeriare e-process. This is due to the following result by Larsson et al. [2024]:
Lemma 6.3 (Theorem 4.1 of Larsson et al. [2024]).
If $Y=\frac{\mathrm{d}Q}{\mathrm{d}P^{*}}(X_{1})$ is an e-value for $\mathcal{P}$ (i.e. $\sup_{P\in\mathcal{P}}\mathrm{E}_{P}(Y)\leqslant 1$) and $P^{*}\in\mathcal{P}$, then, among all e-values for $\mathcal{P}$, $Y$ maximizes $\mathrm{E}_{Q}(\log Y)$ and is called the numeraire e-value.
@@
$P^{*}=N_{0,\sigma^{2}+\mu^{2}}$ attains the following minimum
$\min_{P\in\mathcal{N}_{\mu=0}}D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}\|P)=\frac{1}{2}\log(1+\mu^{2}/\sigma^{2}).$ (65)
thus it is the reverse information projection of $N_{\mu,\sigma^{2}}$ onto $\mathcal{N}_{\mu=0}$. Further, $M_{n}^{*}=\prod_{i=1}^{n}\frac{\mathrm{d}N_{\mu,\sigma^{2}}}{\mathrm{d}N_{0,\sigma^{2}+\mu^{2}}}(X_{i})$ is a supermartingale for $\mathcal{N}_{\mu=0}$, thus the numeraire e-process. Consequently, any e-process $\{M_{n}\}$ for $\mathcal{N}_{\mu=0}$ grows at most
$\frac{\mathrm{E}_{N_{\mu,\sigma^{2}}}(\log M_{n})}{n}\leqslant\frac{1}{2}\log(1+\mu^{2}/\sigma^{2}),$ (66)
equality if and only if $\{M_{n}\}$ is the numeraire $\{M_{n}^{*}\}$.
The proposition above is proved by computing explicitly that $\mathrm{E}_{P_{0}}\left(\frac{\mathrm{d}Q}{\mathrm{d}P^{*}}(X_{1})\right)\leqslant 1$, detailed in Section C.4. It indicates that all test processes summarized in Tables 2 and 3 have optimal growth indices.

### v4

@@ 1 Introduction
Lai and more recent authors [Grünwald et al., 2023, Pérez-Ortiz et al., 2022] situate the Gaussian t-test problem into the broader framework of group-invariant tests; indeed, the t-test null is invariant under rescaling of observations by a constant. The tests are mostly constructed using variants of the sequential likelihood ratio test (LRT) by Wald [1945], utilizing the fact that the likelihood ratios are nonnegative martingales (see Section 2.5). Notably, Pérez-Ortiz et al. [2022] shows the group-invariant extension of the LRT to be optimal (in the “e-power” sense to be defined in Section 2.3) in the orbit-versus-orbit ($\mu/\sigma=\theta_{0}$ versus $\mu/\sigma=\theta_{1}$) setting. However, their results fail to generalize to the composite alternative $\mu/\sigma\neq\theta_{0}$. Building upon these group invariance tools, Lindon et al. [2022] construct a sequential F-test for the regression set-up, which coincides with one of the t-tests that we shall present when no covariate exists. A slightly different approach is taken by Jourdan et al. [2023] who study the same problem in the bandit context, sequentially bounding the empirical mean and variance separately using existing Z-test-type bounds and obtaining via a union bound a sequence of confidence “boxes” for mean and variance simultaneously. We shall return to and compare our bounds with some of these results later.
@@
    We derive the information-theoretic lower bound on the width of confidence intervals, and the upper bound on the growth of e-processes for the t-test, both of which are attained, or almost attained, by our results (Contributions 1 and 3 in this list). We also show that other conservative sequential t-tests available in the literature (Lai’s original method, the plug-in method by Jourdan et al. [2023], the median method by Howard and Ramdas [2022]), and very surprisingly, even the fixed-time classical t-test, fail to be optimal in these regards.
  5. 5.
    We conduct a detailed comparison among tests we derive and in the literature in terms of (a) which filtration they are safe for optional stopping in (the larger the better); (b) what their e-powers are against a fixed alternative (the larger the better); and (c) for their derived confidence sequences, the width or expected width, especially as a function of sample size $n$ and level $\alpha$ (smaller is better).
@@ 2.2 Sequential Statistics
2.3 Test processes and their e-power
@@
In particular, when only $n=1$ observation is considered, we call a nonnegative random variable $M=M(X_{1})$ with $\sup_{P\in\mathcal{P}}\mathrm{E}_{P}(M)\leqslant 1$ an e-value for $\mathcal{P}$.
@@
Lemmas 2.1 and 2.2 suggest that, if $\mathcal{P}_{0}$ is a set of null distributions, NSMs, e-processes, and ENSMs for $\mathcal{P}_{0}$ can be seen as measurements of evidence against the null $H_{0}:P\in\mathcal{P}_{0}$. To see that, if we have at hand an NSM or e-process $\{M_{n}\}$ for $\mathcal{P}_{0}$ such that $M_{0}=1$, we may reject the null $\mathcal{P}_{0}$ whenever the process $M_{n}$ exceeds $1/\alpha$. This, due to Lemma 2.1, controls the sequential type 1 error $P(H_{0}\text{ is ever rejected})$ within $\alpha$ for any $P\in\mathcal{P}_{0}$. Similarly, if we have an ENSM $\{M_{n}\}$ for $\mathcal{P}_{0}$, we may set the right hand side of (5) to $\alpha$ and solve the corresponding $\varepsilon$, rejecting $H_{0}:P\in\mathcal{P}_{0}$ whenver $M_{n}$ exceeds $\varepsilon$. We refer to NSMs, e-processes, and ENSMs for $\mathcal{P}_{0}$ as test processes for the null $\mathcal{P}_{0}$. Desirably, a powerful test process shrinks under the null $\mathcal{P}_{0}$, and grows under the alternative $\mathcal{P}_{1}$ so rejection can happen. To measure the power of a test process under a fixed alternative distribution $Q\in\mathcal{P}_{1}$, we define the e-power of a test process $\{M_{n}\}$ as the $Q$-almost sure limit $\liminf_{n\to\infty}\frac{\log M_{n}}{n}$; for a single e-value $M=M(X_{1})$, its e-power is defined as the expected value $\mathrm{E}_{Q}(\log M)$ [Vovk and Wang, 2024].
A confidence sequence for a parameter $\theta$ follows in the same way if we have an NSM or an ENSM $\{M^{\theta_{0}}_{n}\}$ for every $\mathcal{P}_{\theta_{0}}=\{P\in\mathcal{P}:\theta(P)=\theta_{0}\}$: by solving the inequality $M_{n}^{\theta_{0}}\leqslant 1/\alpha$ or $M_{n}^{\theta_{0}}\leqslant\varepsilon$ in $\theta_{0}$, that is. An illustrative comparison between NSMs and ENSMs in the Gaussian case with known variance for both constructing tests and confidence sequences is provided recently by Wang and Ramdas [2023a, Section 5].
@@ 2.5 Likelihood Ratio Martingales
As a test process, the e-process $\{R_{n}\}$ distinguishes the null $\mathcal{N}_{\mu=0}$ and the alternative $\mathcal{N}_{\mu\neq 0}$ due to the following asymptotic result on a limit which, we recall, is defined in Section 2.3 as the “e-power” of the test process.
Proposition 3.3 (e-power of the universal inference t-test e-process).
@@
The $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ e-power in Proposition 3.3, we remark, is universal among test processes for t-tests. If one uses the plug-in likelihood ratio $\{\ell_{n}^{0,\sigma^{2}}\}$ in Corollary 3.1 to conduct the Z-test for the null $N_{0,\sigma^{2}}$, then, under the actual distribution $N_{\mu,\sigma^{2}}$, it is not hard to see that a faster rate $\frac{\log\ell_{n}^{0,\sigma^{2}}}{n}\to\mu^{2}/2\sigma^{2}$ holds under similar assumptions of the point estimators (which we show in Appendix B). We shall see in the rest of the paper many more occurrences of the same $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ limit, and we shall discuss that this is indeed the optimal rate in Section 6.3.
@@
Proposition 3.5 (e-power of the universal inference one-sided t-test e-process).
@@ 4.4 An Extended Test Martingale for t-Test
$\{H_{n}\}$ works favorably as a test process for the null $\mathcal{N}_{\mu=0}$ and the alternative $\mathcal{N}_{\mu\neq 0}$ due to the following symptotic result on both its e-power under the alternative and its convergence under the null.
@@ 5.1 Theoretical Comparison
e-power $\mu^{2}/2\sigma^{2}$ $(\mu\vee 0)^{2}/2\sigma^{2}$ $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ $\frac{1}{2}\log(1+(\mu\vee 0)^{2}/\sigma^{2})$
Table 2: Comparison of universal inference-based Z-tests and t-tests, where $K_{n}:=\prod_{i=1}^{n}\frac{\exp\left\{-\frac{1}{2}\left(\frac{X_{i}-\widetilde{\mu}_{i-1}}{\widetilde{\sigma}_{i-1}}\right)^{2}\right\}}{\widetilde{\sigma}_{i-1}}$. The “e-power” of a process $M_{n}$ refers to the almost sure limit of $\frac{\log M_{n}}{n}$ under the true distribution $N_{\mu,\sigma^{2}}$.
@@
e-power $\mu^{2}/2\sigma^{2}$ $(\mu\vee 0)^{2}/2\sigma^{2}$ $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ $\frac{1}{2}\log(1+(\mu\vee 0)^{2}/\sigma^{2})$
@@ 6 Optimality
In this section, we show that the $\mathcal{O}(\alpha^{-1/n})$ or $\mathcal{O}(\alpha^{-1/(n-1)})$ growth rate (when $n$ is fixed and $\alpha\to 0$) in our t-confidence sequences shown in Table 1, and that the $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ e-power in our t-test processes shown in Tables 2 and 3, are both optimal. These two kinds of optimality require different techniques, but the Kullback-Leibler divergence
@@ 6.2 The Surprising $\alpha$-Suboptimality of the Classical t-Test
To see that this is an e-value without invoking the full Theorem 3.2, use the fact that $|x-\mu|\leqslant\sigma\exp((x-\mu)^{2}/2\sigma^{2}-1/2)$ for all $x,\sigma$ to conclude that
@@
$\mathrm{E}_{N_{\mu,\sigma^{2}}}\left(\sigma\exp\left(\frac{(X_{1}-\mu)^{2}}{2\sigma^{2}}-\frac{1}{2}\right)\cdot\exp\left(\frac{1-X_{1}^{2}}{2}\right)\right)=\mathrm{E}_{N_{\mu,\sigma^{2}}}\left(\frac{\mathrm{d}N_{0,1}}{\mathrm{d}N_{\mu,\sigma^{2}}}(X_{1})\right)=1.$ (59)
@@
Remark 6.3 (On one-observation e-values).
E-values for $\mathcal{N}_{\mu}$ via one single observation $X_{1}$ can be constructed by methods other than (57). For example, in anticipation of the reverse information projection to be introduced in the next subsection, and fully formulated in Section C.4 by Proposition C.3, the following likelihood ratio,
$B:=\frac{\mathrm{d}N_{0,1}}{\mathrm{d}N_{\mu,\mu^{2}+1}}(X_{1})=\sqrt{\mu^{2}+1}\exp\left(\frac{(X_{1}-\mu)^{2}}{2(\mu^{2}+1)}-\frac{X_{1}^{2}}{2}\right),$ (62)
is an e-value for the entire set $\mathcal{N}_{\mu}$, not only for $N_{\mu,\mu^{2}+1}$, meaning that $\mathrm{E}_{P}[B]\leqslant 1$ for any $P\in\mathcal{N}_{\mu}$. (62), however, does not lead to a closed-form CI via inversion.
Remark 6.4 (Related work on one-observation CIs).
(60) is an instance of “one-observation confidence intervals” for $\mu$ over $\mathcal{N}$, a classical topic studied by numerous authors including Charles M. Stein, Abbott and Rosenblatt [1962], and Portnoy [2019] who studies the lower bound in this problem [^2] The paper by Portnoy [2019] contains an important typo: the two $\Phi(\lambda(1+\tfrac{1}{c_{1}}))$ terms should be $\Phi(\lambda(1-\tfrac{1}{c_{1}}))$ instead, in the main result Theorem 1.1.. These constructions typically look like $[X_{1}\pm c_{\alpha}|X_{1}|]$, where, for example, $c_{0.05}\approx 10$, and are often tighter than (60). Our CI (60), however, enjoys a simple closed-form expression that requires no numerical inversion, and is proved directly by the e-value (57) to provide intuition. All these CIs including our (60), we note, contain 0 (or another fixed constant by shifting) regardless of the value of $X_{1}$. These prior results, together with our (60), showcase the lesser-known suboptimality of the classical t-test CI under small samples; in particular, its unnecessary vacuity when $n=1$. On a separate note, our one-observation CI (60) can be randomly and losslessly tightened due to a recent randomized Markov’s inequality by Ramdas and Manole [2023, Theorem 1.2]. This randomization halves the expected width of the CI. To elaborate, the randomized CI has the $1/\alpha$ term in (60) replaced by a $U/\alpha$, where $U$ is an independent uniform random variable on $[0,1]$, which halves the expected width without changing its validity.
6.3 Log-Optimality by Reverse Information Projection
We now show that the recurring e-power $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ in various test processes presented thus far is not a coincidence, as it is optimal in the problem of t-test. We study the e-power upper bound by focusing, again, on the case with a single observation $X_{1}$, via the expected log-value under a fixed alternative of an e-value (called “e-power” by Vovk and Wang [2024], and “growth rate” [^3] To be disambiguated with the “growth rate” of a confidence interval as $\alpha\to 0$ discussed in our paper. by Pérez-Ortiz et al. [2022]). If an e-value $M=M(X_{1})$ for $\mathcal{P}$ necessarily satisfies the e-power upper bound $\mathrm{E}_{Q}(\log M)\leqslant\delta(Q)$ for $Q\notin\mathcal{P}$, then the e-power of any supermartingale for $\mathcal{P}$ can at most be $\delta(Q)$ under $Q$ as well.
@@ 6.3 Log-Optimality by Reverse Information Projection
It is known that when testing the point null $P$, if the real distribution $Q\neq P$, the e-power of an e-value $M$ is at most
$\mathrm{E}_{Q}(\log M)\leqslant D_{\operatorname{KL}}(Q\|P),$ (63)
@@
with equality if and only if $M$ is the likelihood ratio $\frac{\mathrm{d}Q}{\mathrm{d}P}(X_{1})$, i.e. the $L_{1}$ in Lemma 2.3. See e.g. Vovk and Wang [2024, Lemma 1]. For example, in the Z-test, $P=N_{0,\sigma^{2}}$ and $Q=N_{\mu,\sigma^{2}}$, then the upper bound reads
$D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}\|N_{0,\sigma^{2}})=\frac{\mu^{2}}{2\sigma^{2}},$ (64)
matching those Z-test grow rates in Tables 2 and 3.
@@
As a direct consequence, if we are to test a composite null $\mathcal{P}$ and the real distribution $Q\notin\mathcal{P}$, an e-value $M$ has e-power at most
$\mathrm{E}_{Q}(\log M)\leqslant\inf_{P\in\mathcal{P}}D_{\operatorname{KL}}(Q\|P).$ (65)
With the t-test where $\mathcal{P}=\mathcal{N}_{\mu=0}$ and $Q=N_{\mu,\sigma^{2}}$, a simple calculation shows the following.
@@
The probability measure $N_{0,\sigma^{2}+\mu^{2}}$ attains the following minimum
$D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}\|N_{0,\sigma^{2}+\mu^{2}})=\min_{P\in\mathcal{N}_{\mu=0}}D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}\|P)=\frac{1}{2}\log(1+\mu^{2}/\sigma^{2}).$ (66)
Consequently, the e-power of a t-test e-value is at most $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$ under $N_{\mu,\sigma^{2}}$.
This upper bound is matched by the e-powers of all two-sided t-test processes we derive throughout the paper, as displayed in Tables 2 and 3. We shall prove an extended version of Proposition 6.5 in Section C.4, as Proposition C.3, where we discuss the minimization (66) as an instance of “reverse information projection”, a topic studied recently by Larsson et al. [2024] regarding the design of log-optimal e-values.